OpenStudy (itrymath):

Here is 1 and 1 more

OpenStudy (itrymath):

@jim_thompson5910

OpenStudy (itrymath):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get with this one?

OpenStudy (itrymath):

so the second one i believe is B and the first one A

jimthompson5910 (jim_thompson5910):

how did you get A for the first one?

OpenStudy (itrymath):

none of them make sense tbh

jimthompson5910 (jim_thompson5910):

are you able to find the slope of the line through any two points?

OpenStudy (itrymath):

hold on

jimthompson5910 (jim_thompson5910):

hopefully \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\] seems familiar?

OpenStudy (itrymath):

yes, its 4

jimthompson5910 (jim_thompson5910):

good, so 4 will be in front of the n ie you'll have 4n and not 5n

OpenStudy (itrymath):

so C

jimthompson5910 (jim_thompson5910):

the answer is either 4n+1 or 4n-1 to figure out which it is, plug in any point you want and see which points are satisfied

jimthompson5910 (jim_thompson5910):

yes 4n+1

OpenStudy (itrymath):

+

OpenStudy (itrymath):

YAY omg i didnt think of slope formula thanks alot!!! my finals are next week on Wednesday and i need all help i can get thanks but these are last 2 and then i have 1 more that i know answer too but not sure how to get it

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (itrymath):

the second one is in the picture i already provided you with xD

jimthompson5910 (jim_thompson5910):

any ideas for that one? or are you stuck completely and haven't gotten started?

OpenStudy (itrymath):

no ideas @jim_thompson5910

jimthompson5910 (jim_thompson5910):

have you heard of the Fibonacci sequence?

OpenStudy (itrymath):

Yes a little

OpenStudy (itrymath):

only the name tbh

jimthompson5910 (jim_thompson5910):

the sequence is generated by adding the previous two terms eg: 2+3 = 5 3+5 = 8 5+8 = 13 etc

jimthompson5910 (jim_thompson5910):

in terms of fancy notation, this means f(n) = f(n-1) + f(n-2)

OpenStudy (itrymath):

okay @jim_thompson5910

OpenStudy (itrymath):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

where are you stuck? there's only one answer choice that fits that format

OpenStudy (itrymath):

yeah thats the thing i dont understand

OpenStudy (itrymath):

the second one

jimthompson5910 (jim_thompson5910):

f(n) is the nth term f(n-1) is the term just before the nth term f(n-2) is the term just before the f(n-1) term

jimthompson5910 (jim_thompson5910):

adding the previous terms before the nth term means we add up f(n-1) and f(n-2) to get f(n) = f(n-1) + f(n-2)

jimthompson5910 (jim_thompson5910):

which is exactly how the Fibonacci sequence is defined

OpenStudy (itrymath):

like thats it or is there more??

OpenStudy (itrymath):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's all there is to it. As long as n > 2, the sequence holds up

OpenStudy (itrymath):

what if it was the other way around n was less than 2? @

OpenStudy (itrymath):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

it wouldn't work because n = 1 leads to f(n-2) = f(1-2) = f(-1) which doesn't make much sense f(-1) would refer to the negative first term, which doesn't exist

OpenStudy (itrymath):

okay thanks !! so much i know i tagged you probably like 50 times today annoying the crap out of you but i have 1 more thing!

OpenStudy (itrymath):

do you want new post or this is fine?

jimthompson5910 (jim_thompson5910):

either is fine

OpenStudy (itrymath):

the answer i picked is correct but i found it on google and not sure how they got that answer

jimthompson5910 (jim_thompson5910):

you are correct

jimthompson5910 (jim_thompson5910):

you can plug in each point into that equation to check y = 4(2)^x 8 = 4(2)^1 ... plug in (x,y) = (1,8) 8 = 4(2) 8 = 8 ... works y = 4(2)^x 64 = 4(2)^4 ... plug in (x,y) = (4,64) 64 = 4(16) 64 = 64 ... works

OpenStudy (itrymath):

wow that simple thanks !!!

OpenStudy (itrymath):

wish i could givve more medals :)

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking now

OpenStudy (itrymath):

okay thats it for today will you be on tomorrow ?

jimthompson5910 (jim_thompson5910):

I'm not sure

OpenStudy (itrymath):

well okay thanks bye!!!!

jimthompson5910 (jim_thompson5910):

you're welcome and have a good night